On a farm, there are only sheep and chicken. There were 60 heads and 200 legs. How many sheep are there?
Answer by mananth(16949)

Let sheep be x
and chicken be y
Each has only 1 head
so x+y =60..................(1)
Sheep has 4 legs and chicken 2
4x+2y=200....................(2)
Multiply (1) by -2 and add the equation to (2)
-2x-2y=-120
4x+2y=200
Therefore 2x= 80
/2
x= 40
sheep 40 numbers balance 20 are chickens
CHECK
40*4 +20*2= 200